# SafeInt: Shielding Large Language Models from Jailbreak Attacks via Safety-Aware Representation Intervention

### v1

@@ Abstract
With the widespread real-world deployment of large language models (LLMs), ensuring their behavior complies with safety standards has become crucial. Jailbreak attacks exploit vulnerabilities in LLMs to induce undesirable behavior, posing a significant threat to LLM safety. Previous defenses often fail to achieve both effectiveness and efficiency simultaneously. Defenses from a representation perspective offer new insights, but existing interventions cannot dynamically adjust representations based on the harmfulness of the queries. To address this limitation while ensuring both effectiveness and efficiency, we propose SafeIntervention (SafeInt), a novel defense method that shields LLMs from jailbreak attacks through safety-aware representation intervention. SafeInt is built on our analysis of the representations of jailbreak samples. It adjusts representation distributions of jailbreak samples through intervention to align them with the representations of unsafe samples while minimizing unnecessary perturbations to jailbreak-irrelevant representations. We conduct comprehensive experiments covering six jailbreak attacks, two jailbreak datasets, and two utility benchmarks. Experimental results demonstrate that SafeInt outperforms all baselines in defending LLMs against jailbreak attacks while largely maintaining utility. Additionally, we evaluate SafeInt against adaptive attacks and verify its effectiveness in mitigating real-time attacks.
@@ 1 Introduction
Large Language Models (LLMs) OpenAI et al. (2024); Touvron et al. (2023); Grattafiori et al. (2024) have demonstrated remarkable performance across various domains Zhang et al. (2023); Liu et al. (2023); Wang et al. (2024). With their widespread application in real-world scenarios, LLMs face safety challenges Ferrara (2023); Ji et al. (2023). Although efforts Wang et al. (2023a); Rafailov et al. (2024); Zhou et al. (2023) have been made to align LLMs’ behaviors with human values through carefully designed training strategies, recent studies Zou et al. (2023b); Li et al. (2024); Mehrotra et al. (2024); Liu et al. (2024b) reveal that LLMs can still produce undesirable behaviors when subjected to well-crafted jailbreak attacks, such as the biased generation or potentially harmful responses.
Various defense methods have been proposed to address the growing threat of jailbreak attacks. Prompt-based defenses use instructions Phute et al. (2024); Xie et al. (2023); Zhang et al. (2024) or context Zhou et al. (2024a); Wei et al. (2024) to prevent LLMs from generating harmful content. However, prompt-based methods rely on manually crafted secure prompts and possibly lead to excessive self-censorship Varshney et al. (2024), reducing the helpfulness of LLMs for benign queries. Detection-based defenses compute the perplexity of inputs Alon and Kamfonas (2023) or perturb them Cao et al. (2024) to identify jailbreak prompts. Decoding-based defenses Xu et al. (2024); Liu et al. (2024a) reconstruct a safer output probability distribution through contrastive decoding. However, these methods often lack effectiveness or require additional inference overhead. We aim to defend LLMs against jailbreak attacks from a representation perspective, which provides a more controllable and efficient approach. Previous studies Zou et al. (2023a); Rimsky et al. (2024) have shown the effectiveness of intervening representations to steer LLMs’ behaviors, but such interventions cannot dynamically adjust representations based on whether a query is harmful. This limitation makes it challenging to leverage representations for mitigating jailbreak attacks.
@@
Building on these observations, we propose SafeIntervention (SafeInt), a novel defense method that shields LLMs from jailbreak attacks via safety-aware representation intervention. Since the representations of jailbreak samples differ from those of unsafe samples, LLMs fail to reject jailbreak samples. The core idea of SafeInt is to adjust representation distributions of jailbreak samples through intervention and align them with unsafe representations while minimizing unnecessary perturbations to jailbreak-irrelevant representations. To achieve this, SafeInt adopts a parameterized intervention at an intermediate layer and performs representation alignment in subsequent layers by incorporating classification probabilities and contrastive learning. For jailbreak-irrelevant representations, we perform representation reconstruction to maintain their stability. After training, SafeInt can adaptively identify and intervene in jailbreak-related representations while seamlessly integrating into the LLM inference process.
@@ 2.2 Analysis of Jailbreak Sample Representations
We reconstruct a test set where jailbreak samples are composed of three methods: GCG, AutoDAN Liu et al. (2024b), and DeepInception Li et al. (2024). We employ the previously trained classifiers for testing and show the results in Figure 2. Since the classifiers are trained solely on GCG jailbreak samples, a high test accuracy reveals that the representations generated by different jailbreak methods exhibit a unified pattern from the classifier’s perspective, indicating consistency.
@@
Key Insights and Motivation Aligned LLMs can reject unsafe instructions, but due to the distinction in representations between jailbreak samples and unsafe samples, LLMs fail to reject jailbreak instructions, resulting in the generation of harmful content. Based on the distinguishability of jailbreak sample representations and their observed distributional consistency, we aim to intervene in these representations to align their distribution with that of unsafe samples. This alignment enables the LLM to recognize jailbreak samples as unsafe and reject them accordingly.
@@ 3 Method
3.1 Intervention Parameterization
We aim to find an intervention that, when applied to the representation of the last token in a jailbreak sample, aligns its representation with that of unsafe samples. At the same time, the intervention must not affect the normal responses of safe samples. To this end, we parameterize the intervention. Specifically, we use LoReFT Wu et al. (2024) to parameterize the intervention applied to the last token at a specific layer. LoReFT is an efficient representation fine-tuning method that intervenes on representations within a linear subspace defined by a low-rank projection matrix. Assuming the intervention is applied at layer $\mathcal{I}$, it can be parameterized as follows:
$\widetilde{\mathbf{h}}^{(\mathcal{I})}=\mathbf{h}^{(\mathcal{I})}+\mathbf{U}^{\top}(\mathbf{Wh^{(\mathcal{I})}}+\mathbf{b}-\mathbf{Uh^{(\mathcal{I})}}).$ (2)
The matrix $\mathbf{U}\in\mathbb{R}^{r\times d}$ has orthonormal rows, where $r$ denotes the rank of the subspace, and $\mathbf{W}\in\mathbb{R}^{r\times d}$ and $\mathbf{b}\in\mathbb{R}^{r}$ are the projection matrix and bias vector, respectively.
Then, we define the objectives for learning the intervention. Broadly, our objectives are twofold: a safety objective and a utility objective. The safety objective guarantees the intervention to help the LLM reject jailbreak and harmful instructions. The utility objective ensures that the intervention does not degrade the response quality for harmless instructions, maintaining the LLM’s high availability.
@@ 3.2 Representation Alignment
We first use the classifier as a proxy to assess whether the distributions of jailbreak samples and unsafe samples are consistent in the representation space. From the perspective of the classifier, the alignment is achieved when the classification results for jailbreak and unsafe sample representations are consistent. Specifically, for jailbreak samples, we intervene on their representations to maximize the probability of being classified as unsafe. For unsafe sample representations, they should still be classified as unsafe with a high probability.
@@
Contrastive Learning Although we align the representations of jailbreak samples with unsafe samples by a classifier, the limited training data may prevent the classifier’s decision boundary from accurately capturing the discriminative boundary within the LLM. To enhance the alignment, we use contrastive learning as a complementary task. Contrastive learning is a widely used technique that pulls similar samples closer while pushing different samples apart in the representation space Sohn (2016). For a given representation $\mathbf{q}$, contrastive learning maximizes the similarity between $\mathbf{q}$ and the set of positive samples $K^{+}$ while minimizing the similarity between $\mathbf{q}$ and the set of negative samples $K^{-}$, with the objective formulated as follows:
@@ 3.3 Representation Reconstruction
To prevent excessive intervention from distorting the LLM’s internal representations, we introduce a reconstruction loss to constrain jailbreak-irrelevant representations from changing. Specifically, we encourage the representations of safe and unsafe samples after intervention to remain close to their original states. This ensures that the intervention primarily affects jailbreak-related representations without causing unnecessary shifts in the model’s overall representation structure. The loss is formulated as follows:
@@ 4.1 Experimental Setup
Jailbreak Attacks Multiple representative jailbreak attacks are employed in our evaluation. These include optimization-based attacks: GCG and AutoDAN, LLM-generated attacks: PAIR Chao et al. (2024b) and TAP Mehrotra et al. (2024), and scenario-based attacks: DeepInception. We also consider multilingual mismatch generalization attacks (MG) Yong et al. (2024), where each instruction in the test set is translated into one of six non-English languages to perform the attacks. Detailed setup can be found in Appendix A.2.
@@
Implementation Details Previous classification results indicate that in the intermediate layers, the representations of various jailbreak samples are relatively consistent, and jailbreak samples can be effectively distinguished from both unsafe and safe samples. To avoid time-consuming searches, we directly select layer $\mathcal{I}=12$ as the intervention layer. Since Vicuna exhibits weaker security, deeper alignment is required, so we set its alignment layers from layer $15$ to the final layer. For Llama2, which has undergone safety reinforcement, aligning only the final layer is sufficient.
@@ 4.2 Main Results
Due to the efficiency of LoReFT, SafeInt can be seamlessly integrated into LLM inference, similar to LoRA Hu et al. (2021). Unlike SafeDecoding, which requires an additional expert model for contrastive decoding, SafeInt introduces virtually no extra computational overhead and is therefore more efficient.
@@ 4.3 Adaptive Attack
We also consider a scenario where the attacker knows SafeInt and has access to the LLM deployed with it. This means the attacker can dynamically adjust their attack strategies based on the latest defended LLM to bypass the defense mechanisms more effectively. Table 4 presents the experimental results in this scenario, showing that SafeInt maintains strong defensive performance. After SafeInt is deployed, even if GCG and AutoDAN optimize their adversarial prompts in real time, it is difficult to generate threatening attacks.
@@ 6 Related Work
Jailbreak Defenses To address the challenges posed by jailbreak attacks, numerous defense methods have been proposed Robey et al. (2024); Kumar et al. (2025). Detection-based approaches identify adversarial prompts by computing perplexity Alon and Kamfonas (2023) or randomly deleting parts of the input Cao et al. (2024). Some methods prompt the LLM to perform self-checking through instructions Phute et al. (2024); Xie et al. (2023); Zhang et al. (2024) or context Zhou et al. (2024a). Decoding-based defenses Xu et al. (2024); Liu et al. (2024a) focus on analyzing decoding probabilities under different conditions and formulating decoding strategies to ensure safer outputs. Additionally, certain approaches Zhao et al. (2024); Ouyang et al. (2025) edit specific model parameters to make LLMs forget harmful knowledge. A more controllable and efficient class of defenses Li et al. (2025); Shen et al. (2025) involves manipulating representations to mitigate jailbreak attacks without modifying model parameters or adding decoding overhead.

### v2

@@ Abstract
With the widespread real-world deployment of large language models (LLMs), ensuring their behavior complies with safety standards has become crucial. Jailbreak attacks exploit vulnerabilities in LLMs to induce undesirable behavior, posing a significant threat to LLM safety. Previous defenses often fail to achieve both effectiveness and efficiency simultaneously. Defenses from a representation perspective offer new insights, but existing interventions cannot dynamically adjust representations based on the harmfulness of the queries. To address this limitation, we propose SafeIntervention (SafeInt), a novel defense method that shields LLMs from jailbreak attacks through safety-aware representation intervention. Built on our analysis of the representations of jailbreak samples, the core idea of SafeInt is to relocate jailbreak-related representations into the rejection region. This is achieved by intervening in the representation distributions of jailbreak samples to align them with those of unsafe samples. We conduct comprehensive experiments covering six jailbreak attacks, two jailbreak datasets, and two utility benchmarks. Experimental results demonstrate that SafeInt outperforms all baselines in defending LLMs against jailbreak attacks while largely maintaining utility. Additionally, we evaluate SafeInt against adaptive attacks and verify its effectiveness in mitigating real-time attacks. WARNING: This paper may contain content that is offensive and harmful.
@@ 1 Introduction
Large Language Models (LLMs) OpenAI et al. (2024); Touvron et al. (2023); Grattafiori et al. (2024) have demonstrated remarkable performance across various domains Zhang et al. (2023); Liu et al. (2023); Wang et al. (2024). With their widespread application in real-world scenarios, LLMs face safety challenges Ferrara (2023); Ji et al. (2023). Although efforts Wang et al. (2023a); Rafailov et al. (2024); Zhou et al. (2023) have been made to align LLMs’ behaviors with human values through carefully designed training strategies, recent studies Zou et al. (2023b); Li et al. (2024b); Mehrotra et al. (2024); Liu et al. (2024b) reveal that LLMs can still produce undesirable behaviors when subjected to well-crafted jailbreak attacks, such as the biased generation or potentially harmful responses.
Various defense methods have been proposed to address the growing threat of jailbreak attacks. Prompt-based defenses use instructions Phute et al. (2024); Xie et al. (2023); Zhang et al. (2024b) or context Zhou et al. (2024a); Wei et al. (2024) to prevent LLMs from generating harmful content. However, prompt-based methods rely on manually crafted secure prompts and possibly lead to excessive self-censorship Varshney et al. (2024), reducing the helpfulness of LLMs for benign queries. Detection-based defenses compute the perplexity of inputs Alon and Kamfonas (2023) or perturb them Cao et al. (2024) to identify jailbreak prompts. Decoding-based defenses Xu et al. (2024); Liu et al. (2024a) reconstruct a safer output probability distribution through contrastive decoding. However, these methods often lack effectiveness or require additional inference overhead. We aim to defend LLMs against jailbreak attacks from a representation perspective, which provides a more controllable and efficient approach. Previous studies Zou et al. (2023a); Rimsky et al. (2024) have shown the effectiveness of intervening representations to steer LLMs’ behaviors, but such interventions cannot dynamically adjust representations based on whether a query is harmful. This limitation makes it challenging to leverage representations for mitigating jailbreak attacks.
@@
Building on these observations, we propose SafeIntervention (SafeInt), a novel defense method that shields LLMs from jailbreak attacks via safety-aware representation intervention. The representations of unsafe samples inherently characterize the rejection region of the LLM. However, jailbreak samples often produce representations that deviate from those of unsafe samples, causing the model to fail to trigger its built-in rejection behavior. The core idea of SafeInt is to relocate jailbreak-related representations into the rejection region, thereby activating the model’s native refusal mechanisms. To achieve this, we first project the representations at an intermediate layer into a linear subspace, followed by a parameterized intervention. For jailbreak-related representations, we align their distribution with that of unsafe samples across the subsequent layers. For jailbreak-irrelevant representations, we perform representation reconstruction to preserve their original semantics. After training, SafeInt can adaptively intervene in jailbreak-related representations while seamlessly integrating into the LLM inference process.
@@ 2.2 Analysis of Jailbreak Sample Representations
We reconstruct a test set where jailbreak samples are composed of three methods: GCG, AutoDAN Liu et al. (2024b), and DeepInception Li et al. (2024b). We employ the previously trained classifiers for testing and show the results in Figure 2. Since the classifiers are trained solely on GCG jailbreak samples, a high test accuracy reveals that the representations generated by different jailbreak methods exhibit a unified pattern from the classifier’s perspective, indicating consistency.
@@
Key Insights and Motivation Aligned LLMs can reject unsafe samples, and the representations of these samples inherently characterize the model’s rejection region. Since the representations of jailbreak samples differ from those of unsafe samples, they need to be relocated into the rejection region. Based on the distinguishability and distributional consistency of jailbreak representations, we aim to intervene in their representations to align their distribution with that of unsafe samples.
@@ 3 Method
3.1 Representation Relocation
We achieve representation relocation by a targeted intervention that maps jailbreak-related representations into the rejection region defined by unsafe samples. According to the linear interpretability hypothesis commonly used in existing methods Zhang et al. (2024a); Li et al. (2024a), deep model embeddings can be linearly transformed to correspond to specific human concepts. Thus, we aim to apply a parameterized intervention within a representation space that corresponds to safety-relevant concepts, minimizing impacts on other capabilities. Inspired by LoReFT Wu et al. (2024), we project the representations into a linear subspace defined by a low-rank projection matrix. Assuming the intervention is applied at layer $\mathcal{I}$, it can be parameterized as follows:
$\widetilde{\mathbf{h}}^{(\mathcal{I})}=\mathbf{h}^{(\mathcal{I})}+\mathbf{U}^{\top}\left(f_{\boldsymbol{\theta}}(\mathbf{h^{(\mathcal{I})}})-\mathbf{Uh^{(\mathcal{I})}}\right).$ (2)
The matrix $\mathbf{U}\in\mathbb{R}^{r\times d}$ has orthonormal rows, where $r$ denotes the rank of the subspace. The function $f_{\boldsymbol{\theta}}$ is a linear relocation mapping defined as $f_{\boldsymbol{\theta}}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{r}$.
Then, we define the objectives for learning the intervention. Broadly, our objectives are twofold: a safety objective and a utility objective. The safety objective guarantees the intervention to help the LLM reject jailbreak and harmful instructions. The utility objective ensures that the intervention does not degrade the response quality for harmless instructions.
@@ 3.2 Representation Alignment
We use the classifier as a proxy to assess whether the distributions of jailbreak samples and unsafe samples are consistent in the representation space. From the perspective of the classifier, the alignment is achieved when the classification results for jailbreak and unsafe sample representations are consistent. Specifically, for jailbreak samples, we intervene on their representations to maximize the probability of being classified as unsafe. For unsafe sample representations, they should still be classified as unsafe with a high probability.
@@
Contrastive Learning Although we align the representations of jailbreak samples with unsafe samples by a classifier, the limited training data may prevent the classifier’s decision boundary from accurately capturing the discriminative boundary within the LLM. To enhance the alignment, we use contrastive learning as a complementary task. For a given representation $\mathbf{q}$, contrastive learning maximizes the similarity between $\mathbf{q}$ and the set of positive samples $K^{+}$ while minimizing the similarity between $\mathbf{q}$ and the set of negative samples $K^{-}$, with the objective formulated as follows:
@@ 3.3 Representation Reconstruction
To prevent excessive intervention from distorting the LLM’s internal representations, we introduce a reconstruction loss to constrain jailbreak-irrelevant representations from changing. Specifically, we encourage the representations of safe and unsafe samples after intervention to remain close to their original states. This ensures that the intervention primarily affects jailbreak-related representations. The loss is formulated as follows:
@@ 4.1 Experimental Setup
Jailbreak Attacks Multiple representative jailbreak attacks are employed in our evaluation. These include optimization-based attacks: GCG and AutoDAN, LLM-generated attacks: PAIR Chao et al. (2024b) and TAP Mehrotra et al. (2024), and scenario-based attacks: DeepInception. We also consider multilingual mismatch generalization attacks (MG) Yong et al. (2024), where each instruction in the test set is translated into one of six non-English languages to perform the attacks.
@@
Implementation Details Previous classification results indicate that in the intermediate layers, the representations of various jailbreak samples are relatively consistent, and jailbreak samples can be effectively distinguished from both unsafe and safe samples. To avoid time-consuming searches, we directly select layer $\mathcal{I}=12$ as the intervention layer. Since Vicuna lacks harmless alignment, it exhibits weaker safety. Accordingly, we set the second half of the layers as the alignment layers. In contrast, for models like Llama2 that have undergone safety alignment, aligning only the final layer is sufficient.
@@ 4.2 Main Results
Since our intervention essentially involves an incremental computation, it can be integrated directly into the forward propagation of the model. Unlike SafeDecoding, which requires an additional expert model for contrastive decoding, SafeInt introduces virtually no extra computational overhead and is therefore more efficient.
@@ 4.3 Adaptive Attack
We also consider a scenario where the attacker knows SafeInt and has access to the LLM deployed with it. This means the attacker can dynamically adjust their attack strategies based on the latest defended LLM. The experimental results in Table 4 show that SafeInt maintains strong defense performance under such adaptive settings. Because SafeInt acts inside a low-rank subspace that exists only within the model, no explicit clues about this subspace are exposed at the prompt level or in the externally observable gradients. To bypass SafeInt, an attacker must precisely manipulate specific components of the original representations within this subspace. Therefore, even if GCG and AutoDAN optimize their adversarial prompts in real time, it is difficult to generate threatening attacks.
@@ 6 Related Work
Jailbreak Defenses To address the challenges posed by jailbreak attacks, numerous defense methods have been proposed Robey et al. (2024); Kumar et al. (2025). Detection-based approaches identify adversarial prompts by computing perplexity Alon and Kamfonas (2023) or randomly deleting parts of the input Cao et al. (2024). Some methods prompt the LLM to perform self-checking through instructions Phute et al. (2024); Xie et al. (2023); Zhang et al. (2024b) or context Zhou et al. (2024a). Decoding-based defenses Xu et al. (2024); Liu et al. (2024a) focus on analyzing decoding probabilities under different conditions and formulating decoding strategies to ensure safer outputs. Additionally, certain approaches Zhao et al. (2024); Ouyang et al. (2025) edit specific model parameters to make LLMs forget harmful knowledge. A more controllable and efficient class of defenses Li et al. (2025); Shen et al. (2025) involves manipulating representations to mitigate jailbreak attacks without modifying model parameters or adding decoding overhead.